OpenStudy (anonymous):

Solve.3 over x = 1 over x minus 2

OpenStudy (anonymous):

\[ \frac{3}{x}=\frac{1}{x}-2 \]Is that the question? If so, multiply both sides by x then solve normally.\[ \frac{3}{x}=\frac{1}{x-2} \]If that's the question, cross-multiply.

OpenStudy (anonymous):

I'm confused.

OpenStudy (anonymous):

About what?

OpenStudy (anonymous):

The whole concept

OpenStudy (anonymous):

What whole concept? You're being very vague.

OpenStudy (anonymous):

This problem is confusing to me, i know i need to cross multiply, but i am still not understand this problem.

OpenStudy (anonymous):

So is it the second problem, then? And if so, and you know you need to cross-multiply... what's the problem? Just do the cross-multiplication.

OpenStudy (anonymous):

3/x=1/x-2 multiply by x both sides 3=1-2x subtract 1 from both sides 2=-2x divde both sides by -2 x=-1

OpenStudy (anonymous):

for pandas first case